OpenStudy (anonymous):

solve the equation? I 3x + 2 I =5

OpenStudy (anonymous):

3x + 2 = 5 -3x - 2 = 5 3x = 3 x = 1 -3x = 7 x = -7/3 x = 1, -7/3